# From Stars to Molecules: AI Guided Device-Agnostic Super-Resolution Imaging

### v1

@@ III Conclusion
We demonstrated a device-agnostic approach to super-resolution imaging of point-like emitting sources that utilizes deep learning techniques. The presented method solves the long-standing problems of measuring calibration data and estimating parameters accompanying the training of a deep learning model. Using only numerically simulated data, we develop a model capable of reconstructing intensity images using a single frame without explicit knowledge of the imaging system. This model can be applied to images of arbitrary shape and size originating from various optical systems, and the calibration-free nature grants it resistance against non-uniformity and drifts of the optical parameters. We compared this device-agnostic model with the well-established Richardson-Lucy deconvolution algorithm. The analysis results undeniably show the superior performance of our approach, outperforming the Richardson-Lucy algorithm by orders of magnitude in terms of both reconstruction accuracy and speed. Moreover, we designed an advanced optical setup for acquiring experimental images of emitting sources together with their precise ground truth references. Providing full control over the spatial distribution of emitters, this setup enables exact performance quantification of any reconstruction method. Using these experimental data pairs, we verified the superior performance of our approach.
@@ Acknowledgments
This work was supported by the Czech Science Foundation (project 21-18545S), and the Ministry of Education, Youth, and Sports of the Czech Republic (project OP JAC CZ.02.01.01/00/23_021/0008790). DV acknowledges the support by Palacký University (project IGA-PrF-2024-008).
@@
IV Materials and Methods
IV.1 Data Simulation
The generation process of a single simulated data sample follows these steps. First, we generate the emitter concentration, followed by assigning each emitter its power and pixel position in the image. Next, we perform a convolution with the point spread function to simulate the effects of finite resolution. Subsequently, shot noise is added to each pixel of the image. With knowledge of the average noise intensity and emitter power, we can calculate the signal-to-noise ratio of the generated sample as SNR=emitter poweraverage noise $\text{SNR}=\frac{\text{emitter power}}{\text{average noise intensity}}$. Alternatively, we can calculate the peak-to-noise ratio, PNR=peak emitter intensityaverage noise $\text{PNR}=\frac{\text{peak emitter intensity}}{\text{average noise intensity}}$, using the maximum value of the convolved emitter intensity. It is worth noting that the simulation process can be further enhanced by incorporating additional features and parameters, such as point spread function asymmetry and aberrations [45, 46]. For this showcase study, we opted to simplify the simulation to reduce the computational demand during training. Despite this simplification, the DAMN model achieves excellent results on experimental data and demonstrates significant generalization ability beyond the expected regime, as shown in Fig. 2 (D).
$G(r,\sigma)\sim\text{exp}\left(-\frac{r^{2}}{\sigma^{2}}\right),$
IV.2 Deep Learning Model
IV.3 Richardson-Lucy Algorithm
[…]
IV.4 Experiment
The optical setup, which provides resolution-limited experimental images with their corresponding target objects, consists of three segments, see Fig. 7. The first segment prepares the illumination of the digital micromirror device (DMD). An incoherent broadband light with a central wavelength of 555 nm, produced by a laser-induced fluorescence light source (Crytur MonaLIGHT B01), is coupled to a 4 mm diameter light guide. The emerging light is collimated by a microscope objective $\text{MO}_{\text{s}}$ (OLYMPUS 20x/0.4). An iris diaphragm D is positioned proximal to the objective. This diaphragm is re-imaged onto the DMD (Texas Instruments DLP LightCrafter 6500) by a telescope consisting of two achromatic doublets, $\text{L}_{\text{s1}}$ and $\text{L}_{\text{s2}}$, with focal lengths of 50 mm and 150 mm, respectively. The telescope improves the illumination homogeneity, and re-imaging the diaphragm reduces reflections from the passive parts of the DMD.
In the second setup segment, point-like emitters with the targeted spatial distribution are prepared. By addressing the DMD, we impose a ground truth mask with each pixel comprising a $5\times 5$ micromirror grid with a 7.6 $\mu$m micromirror pitch. As we address only the central micromirror of the grid, the closest distance between two emitting sources is 38 $\mu$m in the DMD plane, see Fig. 7. This mask is re-imaged into the front focal plane of $\text{MO}_{\text{p}}$ by a preparation telescope consisting of a 200 mm focal-length achromatic doublet $\text{L}_{\text{p}}$ and a high-resolution microscope objective $\text{MO}_{\text{p}}$ (OLYMPUS 100x/0.9). As the point-like emitters, of the approximate 310 nm full width at half maximum and 340 nm minimal distance, are created in this plane, we term it the emitter plane.
Figure 7: A detailed visualization of the optical setup. In the illumination path, we generate an incoherent light collimated by a microscopic objective $\text{MO}_{\text{s}}$. The subsequent iris diaphragm D and two achromatic doublets, $\text{L}_{\text{s1}}$ and $\text{L}_{\text{s2}}$, improve the illumination homogeneity and reduce undesired reflections from the passive DMD parts. In the preparation path, the DMD-induced ground truth mask is re-imaged by an achromatic doublet $\text{L}_{\text{p}}$ and a high-resolution microscope objective $\text{MO}_{\text{p}}$, creating point-like emitters in the emitter plane. Lastly, this plane is projected by a low-resolution microscope objective $\text{MO}_{\text{I}}$ onto a camera with a mounted spectral filter.
Unlike the first two segments, which prepare the emitters and provide complete control over their spatial distributions, the final segment represents an imaging setup with optical parameters unknown to the DAMN model. A low-resolution microscope objective $\text{MO}_{\text{I}}$ (OLYMPUS 10x/0.25) projects the emitter plane onto a CMOS camera (ZWO-ASI 178MM, 2.4 $\mu$m square pixels) with a mounted spectral filter (central wavelength of 532 nm with 10 nm spectral width). The $\text{MO}_{\text{I}}$ resolution is approximately 1.13 $\mu$m, regarding the full width at half maximum in the emitter plane. In comparison, the 340 nm minimal emitter distance is $\times3$ below this resolution limit. The emitter plane is imaged with an approximate 230×230$\times$230 × magnification to match the major part of the camera chip. The exposure time for each measurement is set to 5.7 seconds to utilize the majority of the 14-bit dynamic range of the camera while avoiding its saturation. The captured image is down-scaled by a factor of 32 to match the $50\times 50$ ground-truth mask size to enable metric evaluations. Using this imaging setup, we collected the experimental dataset of resolution-limited camera images and their corresponding ground truth masks for various emitter concentrations, see Fig. 4 (A). Additionally, we measured a high-SNR calibration set for precise estimation of the point spread function, which we provided only to the Richardson-Lucy algorithm. Besides the concentration ranging from 50 to 500 emitters, the optical parameters in the collected dataset exhibit the following estimated values: SNR $=2300$, average noise intensity $=10$, and an Airy-shaped kernel with $\sigma=2.05$ px, equivalent to the 3.4 px full width at half maximum. The Rayleigh resolution limit of our imaging system, given by the radius R of the first dark ring in the Airy pattern, is approximately 3.9 px, see Fig. 4 (B).

### v2

@@ Discussion
We demonstrated a device-agnostic approach to super-resolution imaging of point-like emitting sources that utilizes deep learning techniques. The presented approach solves the long-standing problems of measuring calibration data and estimating parameters accompanying the training of a deep learning model. Using only numerically simulated data, we develop a model capable of reconstructing intensity images using a single frame without explicit knowledge of the imaging system. This model can be applied to images of arbitrary shape and size originating from various optical systems, and the calibration-free nature grants it resistance against non-uniformity and drifts of the optical parameters. We compared this device-agnostic model with the well-established Richardson-Lucy deconvolution algorithm. The analysis results undeniably show the superior performance of our approach, outperforming the Richardson-Lucy algorithm by orders of magnitude in terms of both reconstruction accuracy and speed. Moreover, we designed an advanced optical setup for acquiring experimental images of emitting sources together with their precise ground-truth references. Providing full control over the spatial distribution of emitters, this setup enables exact performance quantification of any reconstruction method. Using these experimental data pairs, we verified the superior performance of our approach.
@@
Methods
Data Simulation
The generation process of a single simulated data sample follows these steps. First, we generate the emitter concentration, followed by assigning each emitter its power and pixel position in the image. Next, we perform a convolution with the point spread function to simulate the effects of finite resolution. Subsequently, shot noise is added to each pixel of the image. With knowledge of the average noise intensity and emitter power, we can calculate the signal-to-noise ratio of the generated sample as SNR=emitter poweraverage noise $\text{SNR}=\frac{\text{emitter power}}{\text{average noise intensity}}$. Alternatively, we can calculate the peak-to-noise ratio, PNR=peak emitter intensityaverage noise $\text{PNR}=\frac{\text{peak emitter intensity}}{\text{average noise intensity}}$, using the maximum value of the convolved emitter intensity. It is worth noting that the simulation process can be further enhanced by incorporating additional features and parameters, such as point spread function asymmetry and aberrations [40, 41]. For this showcase study, we opted to simplify the simulation to reduce the computational demand during training. Despite this simplification, the DAMN model achieves excellent results on experimental data and demonstrates significant generalization ability beyond the expected regime, as shown in Fig. 2 d.
$G(r,\sigma)\sim\text{exp}\left(-\frac{r^{2}}{\sigma^{2}}\right),$
Deep Learning Model
Richardson-Lucy Algorithm
[…]
Experiment
The optical setup, which provides resolution-limited experimental images with their corresponding target objects, consists of three segments, see Extended Data Fig. 1. The first segment prepares the illumination of the digital micromirror device (DMD). An incoherent broadband light with a central wavelength of 555 nm, produced by a laser-induced fluorescence light source (Crytur MonaLIGHT B01), is coupled to a 4 mm diameter light guide. The emerging light is collimated by a microscope objective $\text{MO}_{\text{s}}$ (OLYMPUS 20x/0.4). An iris diaphragm D is positioned proximal to the objective. This diaphragm is re-imaged onto the DMD (Texas Instruments DLP LightCrafter 6500) by a telescope consisting of two achromatic doublets, $\text{L}_{\text{s1}}$ and $\text{L}_{\text{s2}}$, with focal lengths of 50 mm and 150 mm, respectively. The telescope improves the illumination homogeneity, and re-imaging the diaphragm reduces reflections from the passive parts of the DMD.
In the second setup segment, point-like emitters with the targeted spatial distribution are prepared. By addressing the DMD, we impose a ground-truth mask with each pixel comprising a $5\times 5$ micromirror grid with a 7.6 $\mu$m micromirror pitch. As we address only the central micromirror of the grid, the closest distance between two emitting sources is 38 $\mu$m in the DMD plane, see Extended Data Fig. 1. This mask is re-imaged into the front focal plane of $\text{MO}_{\text{p}}$ by a preparation telescope consisting of a 200 mm focal-length achromatic doublet $\text{L}_{\text{p}}$ and a high-resolution microscope objective $\text{MO}_{\text{p}}$ (OLYMPUS 100x/0.9). As the point-like emitters, of the approximate 310 nm full width at half maximum and 340 nm minimal distance, are created in this plane, we term it the emitter plane.
Unlike the first two segments, which prepare the emitters and provide complete control over their spatial distributions, the final segment represents an imaging setup with optical parameters unknown to the DAMN model. A low-resolution microscope objective $\text{MO}_{\text{I}}$ (OLYMPUS 10x/0.25) projects the emitter plane onto a CMOS camera (ZWO-ASI 178MM, 2.4 $\mu$m square pixels) with a mounted spectral filter (central wavelength of 532 nm with 10 nm spectral width). The $\text{MO}_{\text{I}}$ resolution is approximately 1.13 $\mu$m, regarding the full width at half maximum in the emitter plane. In comparison, the 340 nm minimal emitter distance is $\times3$ below this resolution limit. The emitter plane is imaged with an approximate 230×230$\times$230 × magnification to match the major part of the camera chip. The exposure time for each measurement is set to 5.7 seconds to utilize the majority of the 14-bit dynamic range of the camera while avoiding its saturation. The captured image is down-scaled by a factor of 32 to match the $50\times 50$ ground-truth mask size to enable metric evaluations. Using this imaging setup, we collected the experimental dataset of resolution-limited camera images and their corresponding ground-truth masks for various emitter concentrations, see Fig. 4 a. Additionally, we measured a high-SNR calibration set for precise estimation of the point spread function, which we provided only to the Richardson-Lucy algorithm. Besides the concentration ranging from 50 to 500 emitters, the optical parameters in the collected dataset exhibit the following estimated values: SNR $=2300$, average noise intensity $=10$, and an Airy-shaped kernel with $\sigma=2.05$ px, equivalent to the 3.4 px full width at half maximum. The Rayleigh resolution limit of our imaging system, given by the radius R of the first dark ring in the Airy pattern, is approximately 3.9 px, see Fig. 4 b.
@@ Acknowledgments
This work was supported by the Czech Science Foundation (project 21-18545S), and the Ministry of Education, Youth, and Sports of the Czech Republic (project OP JAC CZ.02.01.01/00/23_021/0008790). DV acknowledges the support by Palacký University (projects IGA-PrF-2024-008 and IGA-PrF-2025-010).